# Gaussian Process Diffeomorphic Statistical Shape Modelling Out Performs Angle Based Methods for Assessment of Hip Dysplasia

### v1

@@ Abstract
Dysplasia is a recognised risk factor for osteoarthritis (OA) of the hip, early diagnosis of dysplasia is important to provide opportunities for surgical interventions aimed at reducing the risk of hip OA. We have developed a pipeline for semi-automated classification of dysplasia using volumetric CT scans of patients’ hips and a minimal set of clinically annotated landmarks, combining the framework of the Gaussian Process Latent Variable Model with diffeomorphism to create a statistical shape model, which we termed the Gaussian Process Diffeomorphic Statistical Shape Model (GPDSSM). We used 192 CT scans, 100 for model training and 92 for testing. The GPDSSM effectively distinguishes dysplastic samples from controls while also highlighting regions of the underlying surface that show dysplastic variations. As well as improving classification accuracy compared to angle-based methods (AUC $96.2\%$ vs $91.2\%$), the GPDSSM can save time for clinicians by removing the need to manually measure angles and interpreting $2D$ scans for possible markers of dysplasia.
@@
[b]organization=The Ottawa Hosptial, city=Ottawa, country=Canada
@@ 1 Introduction
The current clinical methodology for detecting dysplasia is typically a radiographic or CT-based (2-dimensional) assessment of numerous angular and ratio measurements [6]. Amongst commonly used measures are the Lateral Centre Edge Angle (LCEA) and Acetabular Index (AI). For both measurements, there exist clinically agreed upon thresholds for LCEA (<20°)absent20°(<\$\$)( < 20  ° ) and AI (>15°absent15°>\$\$> 15  °) for identifying dysplasia, as well as thresholds for normality (LCEA>25°absent25°>\$\$> 25  °). Values of LCEA and AI between these ranges are considered borderline and need further evaluation. Given a CT scan of a $2D$ slice of the hip joint, clinicians may measure these angles using visualization software that is designed for radiological investigations. See Figure 1 for how these angles are measured.
Figure 1: Example pelvic radiograph of a dysplastic subject, showing angle measurements, with LCEA angle of 15.7°15.7°15.7  ° and AI of 19.1°19.1°19.1  °.
However, the measured angles on the scan can be sensitive to the patient’s positioning during the scan, pelvic tilt, and the choice of $2D$ slice [7, 8, 9]. Furthermore, this projection necessarily discards valuable information about the underlying $3D$ geometry of the acetabulum. Indeed, recent studies [10, 11] have indicated that ambiguous ‘borderline’ dysplastic cases using the angle-classification system are better resolved by studying angles/ratios from $3D$ CT scans that take into account the full underlying geometry. It remains challenging and time-consuming for clinicians to manually measure multiple angle and ratio quantities on $3D$ scans.
@@ 1.1 Existing Work
There have been multiple recent machine learning approaches [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] in the dysplasia detection literature, aiming to detect dysplasia directly from scans. Many of these works rely on training CNN or UNET architectures on raw $2$D/$3D$ CT and ultrasound scans, to classify/score patients for dysplasia. As with all machine learning approaches, these are very dependent upon the quality and quantity of available training data, in particular deep-learning approaches require substantial amounts of data, some authors quoting a minimum of 2300 training data sets [29]. Much of the published reports utilize relatively small training-data sets.
@@ 1.2 Aim
Our aim was to create a novel semi-automated method based on shape modelling of the acetabulum to identify subjects with clinically relevant dysplasia.
@@ 2.1 Pre-processing to Extract Surface Data
The SSM that we used in this work (see Section 2.3) is trained on surface data. As such, we pre-processed the dataset to extract the regions of interest of the hip surface, in particular the acetabular surface.
The first step was to segment the pelvic region of the scan and extract surface data from the segmented scans. This was performed using the VTK toolkit [38], using a marching-cubes algorithm [39], designed to extract an isosurface from functional input data. For each scan in the dataset, the output of this algorithm produced a triangulated pelvic surface. The obtained surfaces were high resolution, with number of triangles $N\geq 10^{4}$, and triangles are not in geometric correspondence across the dataset; we return to this point in A when we discuss the model fitting process. Next, we extracted the surfaces of the acetabuli. We used Algorithm 1 for extracting this region in a semi-automated manner. This algorithm was applied separately to the left- and right-acetabular regions for each subject; see Figure 2.
Algorithm 1 Cup extraction algorithm
@@
This method is semi-automated; it requires marking of four points on the acetabular rims for the first step. This information was already available in our dataset. This information is relatively fast to obtain for new scan data through manual annotation using visualization software. This step could be automated in future work.
@@
The final pre-processing step involved rigidly aligning all subjects’ left and right acetabula. This removes all sources of variability between patients unrelated to the intrinsic geometric variation of the acetabular surface, namely rotation, scale and translation. This is a standard procedure in the pipeline for fitting a surface model. In the absence of point correspondences, we used a correspondence-less rigid-alignment method using the varifolds shape metric. All patient surfaces were aligned to a randomly chosen reference patient surface. The alignment algorithm is detailed in the Appendix. We show exemplar post-alignment results in Figure 3.
Figure 3: Pre-alignment (top row) and post-alignment (bottom row) for left- and right-acetabulum (in corresponding column).
@@ 2.3 Shape Modelling
The SSMs were fitted using 100 CT scans, consisting of 43 scans of subjects with dysplasia and 57 scans of control subjects.
@@ 2.4 Classification in Latent Space
Once the SSMs and classifiers were trained, it was then possible to present previously unseen data to the SSMs for latent point inference and then use the classifiers to determine the probability of dysplasia. Both of these steps, latent point inference and classification, are computationally inexpensive and fast. This provides a simple and efficient procedure for screening new patients for dysplasia.
@@ 2.5 Testing and Comparison with Angle Based Classification
The testing and comparison were based on 92 CT data sets, which were not used for training of the SSMs and classifiers. These were from 40 control and 52 dysplasia subjects, for these subjects the current clinically accepted LCEA and AI angles were available. Angle-based classification was performed by combining the LCEA and AI angles, using LCEA (<20°)absent20°(<\$\$)( < 20  ° ) and AI (>15°absent15°>\$\$> 15  °) to indicate the presence of dysplasia. The comparison between the angle-based and two SSM-based classifers, namely LDDMM and GPDSSM, performance was made by plotting ROC curves and computing AUC scores.
@@ 3.2 Classification Performance
The GPDSSM outperforms the angle-based representation in terms of AUC ($96.2$% to $91.2$%), suggesting that the overall classification based on our latent-space representation is more discriminative across all classification thresholds in separating dysplastics from controls. The GPDSSM representation also yields an improved sensitivity (performance on dysplastic) over the angle-based representations ($94.2$% to $84.6$%), suggesting that this representation is particularly good at identifying dysplastic cases. In contrast, angle-based representations are more likely to incorrectly classify dysplastic as controls. This is an important area of improvement, as type II errors are far more costly in such a clinical setting, where patients may face significant deterioration of health if not treated due to false negatives.
In Table 1, the GPDSSM outperforms the standard LDDMM across all scores, which indicates the increased expressivity and powerful nonlinear dimensionality reduction afforded by our GPLVM-based diffeomorphic framework. We also remark that while the angle-based method has the best specificity (performance on controls) of $90$%, it underperforms in terms of sensitivity against both LDDMM ($92.3$% to $84.6$%) and GPDSSM ($94.2$% to $84.6$%). This suggests the use of latent-space representations based on the entire three-dimensional geometry has a measurable advantage in detecting dysplastic cases compared to the simpler angle-based representations.
@@ 4 Discussion
In this work, we have developed a semi-automated algorithm to classify the presence of dysplasia from volumetric CT scans, using a novel Gaussian process Statistical Shape Model. We have demonstrated the performance of our method on a dataset of real-world patient scans, our GPDSSM model can reliably detect the presence of dysplasia. Furthermore, our GPDSSM method (accuracy 90.2% and AUC 96.2%) demonstrably improves on predictions based on classical angle measures such as AI and LCEA (accuracy 87.0% and AUC 91.2%), as well as traditional LDDMM-based computational anatomy methods (accuracy 88.0% and AUC 92.7%). Our model utilises the full geometry of the acetabular surface to detect variations most correlated with dysplastic status, and is not dependent on a fixed $2$D CT slice and angle of acquisition/measurement accuracy. The speed of inference and classification algorithm can save valuable time for clinicians, removing the need for manually marking and measuring angles, as well as improving detection accuracy. Given a new patient, clinicians may use our trained GPDSSM to infer dysplastic status as follows.
@@
If the patient is dysplastic, one may further use the type of deformation heat-map analysis presented in Section 2.6, for understanding the mode of disease variation in a patient scan, via the following steps.
@@ A.1 Model specification, fitting and inference
where we further condition on $p(U)$ a prior distribution over inducing variables $U$ [46] in time and latent space, where $U$ is equal to $\alpha$ evaluated at $I_{U}$, a set of $m<N$ free inducing points to be optimized in $[0,1]\times Z$. One defines $p(\bm{\alpha}\mid U,Z)$ through standard Gaussian conditioning formulae.
@@ C.1 Gaussian Process Latent Variable Model
where $y$ is observed data, and $f:Z\longrightarrow Y$ is a randomized mapping from latent space ${Z}$, to ${Y}$ the data space. The noise in the data is typically modelled by a Gaussian random variable. Importantly, $f$ is a Gaussian process [47], meaning the evaluations of $f$ at any finite set of latent points has a joint Gaussian distribution with mean and covariance determined by a mean function $m_{\theta}(z)$ and covariance kernel $k_{\theta}(z,z^{\prime})$ with learnable hyperparameters $\theta$. This is often denoted compactly as
@@ C.4 Classification of shape data
obtained and pre-processed with the methodology presented in Section 2, where $S_{i}$ denote the surface of the $i$’th subject anatomy and $y_{i}\in\{0,1\}$ the corresponding label; $1$ for dysplastics and 00 for controls. After fitting $\mathcal{M}$ to the shapes $S_{i}$, this yields a set

### v2

@@ Abstract
Dysplasia is a recognised risk factor for osteoarthritis (OA) of the hip, early diagnosis of dysplasia is important to provide opportunities for surgical interventions aimed at reducing the risk of hip OA. We have developed a pipeline for semi-automated classification of dysplasia using $3D$ surface models obtained from volumetric CT scans of patients’ hips and a minimal set of four clinically annotated landmarks on the acetabular rim (the most proximal, distal, anterior and posterior aspects), combining the framework of the Gaussian Process Latent Variable Model with diffeomorphism to create a statistical shape model, which we termed the Gaussian Process Diffeomorphic Statistical Shape Model (GPDSSM). We used 192 CT scans, 100 for model training and 92 for testing. The GPDSSM effectively distinguishes dysplastic samples from controls while also highlighting regions of the underlying surface that show dysplastic variations. As well as improving classification accuracy compared to angle-based methods (AUC $96.2\%$ vs $91.2\%$), the GPDSSM can save time for clinicians by removing the need to manually measure angles and interpreting $2D$ scans for possible markers of dysplasia.
@@
[b]organization=The Ottawa Hospital, city=Ottawa, country=Canada
@@ 1 Introduction
The current clinical methodology for detecting dysplasia is typically a radiographic or CT-based (2-dimensional) assessment of numerous angular and ratio measurements [6]. Amongst commonly used measures are the Lateral Centre Edge Angle (LCEA) and Acetabular Index (AI). For both measurements, there exist clinically agreed upon thresholds for LCEA (<20°)(<\$\$) and AI (>15°>\$\$) for identifying dysplasia, as well as thresholds for normality (LCEA>25°>\$\$). Values of LCEA and AI between these ranges are considered borderline and need further evaluation. Given a CT scan of a $2D$ slice of the hip joint, clinicians may measure these angles using visualization software that is designed for radiological investigations. See Figure 1 for how these angles are measured.
Figure 1: Example pelvic radiograph of a dysplastic subject, showing angle measurements, with LCEA angle of 15.7° and AI of 19.1°.
However, the measured angles on the scan can be sensitive to the patient’s positioning during the scan, pelvic tilt, and the choice of $2D$ slice [7, 8, 9]. Furthermore, this projection necessarily discards valuable information about the underlying $3D$ geometry of the acetabulum. Indeed, recent studies [10, 11] have indicated that ambiguous ‘borderline’ dysplastic cases using the angle-classification system are better resolved by studying angles/ratios from $3D$ CT scans that take into account the full underlying geometry. Whilst CT scans are now commonly used for individuals presenting with symptoms consistent with dysplasia, it remains challenging and time-consuming for clinicians to manually measure multiple angle and ratio quantities on $3D$ scans, and often the data is treated like a $2D$ image.
@@ 1.1 Existing Work
There have been multiple recent machine learning approaches [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] in the dysplasia detection literature, aiming to detect dysplasia directly from scans. Many of these works rely on training CNN or UNET architectures on raw 2D2D/$3D$ CT and ultrasound scans, to classify/score patients for dysplasia. As with all machine learning approaches, these are very dependent upon the quality and quantity of available training data, in particular deep-learning approaches require substantial amounts of data, some authors quoting a minimum of 2300 training data sets [29]. Much of the published reports utilize relatively small training-data sets.
@@ 1.2 Aim
Our aim was to create a novel semi-automated method based on shape modelling of the acetabulum to identify subjects with clinically relevant dysplasia. Currently, symptomatic individuals often risk being considered ”normal” based on simple 2D assessments, but on detailed 3D assessments they are not. Thus, unless reviewed by experienced clinicians or have further detailed analysis, these individuals remain symptomatic and untreated for a long time. Improving diagnostic accuracy with new assessments would be very beneficial for this group, allowing them to be diagnosed and offered treatment earlier. The methods used were chosen as they are particularly suited to smaller data sets.
@@ 2.1 Pre-processing to Extract Surface Data
The SSM that we used in this work (see Section 2.3) is trained on surface data. As such, we pre-processed the CT data to extract the regions of interest of the pelvic surface, in particular the acetabular surface.
The first step was to segment the pelvic region of the scan and extract surface data from the segmented scans to create a pelvic surface model. This was performed using the VTK toolkit [38], using a marching-cubes algorithm [39], designed to extract an isosurface from functional input data. For each scan in the dataset, the output of this algorithm produced a triangulated pelvic surface. The obtained surfaces were high resolution, with number of triangles $N\geq 10^{4}$. Next, we extracted the surfaces of the acetabuli. We used Algorithm 1 for extracting this region in a semi-automated manner. This algorithm was applied separately to the left- and right-acetabular regions for each subject; see Figure 2.
Algorithm 1 Acetabulum extraction algorithm
@@
This method is semi-automated; it requires marking, on the pelvic surface model, of four points (the most proximal, distal, anterior and posterior aspects) on the acetabular rims for the first step of initial alignment. This information was already available in our dataset. This information is relatively fast to obtain for new scan data through manual annotation using visualization software. This step could be automated in future work.
@@
The final pre-processing step involved rigidly aligning all subjects’ left and right acetabula. This removes all sources of variability between patients unrelated to the intrinsic geometric variation of the acetabular surface, namely rotation, scale and translation. This is a standard procedure in the pipeline for fitting a surface model. In the absence of point correspondences, we used a correspondence-less rigid-alignment method using the varifolds shape metric [41]. The varifold shape metric measures how different two shapes are by comparing both their geometric surfaces and the directions those surfaces face, making it ideal for quantifying differences between bone structures without needing to match specific points between them. Think of it as a sophisticated ruler that accounts for both the “bumpiness” and “tilt” of bone surfaces, giving you a single number that represents how much one bone shape deviates from another. All patient surfaces were aligned to a randomly chosen reference patient surface. The alignment algorithm is detailed in the Appendix. We show exemplar post-alignment results in Figure 3.
Figure 3: Pre-alignment (top row) and post-alignment (bottom row) for left- and right-acetabulum (in corresponding column), the the left and right surfaces were aligned individually using the green surface as the master surface for alignment.
@@ 2.3 Shape Modelling
The SSMs were fitted using the surface data from 100 CT scans, consisting of 43 scans of subjects with dysplasia and 57 scans of control subjects.
@@ 2.4 Classification in Latent Space
Once the SSMs and classifiers were trained these were then used to determine the probability of dysplasia in the test data set, Figure 4 shows the overall process of model fitting and usage.
Pre-process Fit statistical shape model Classify as dysplastic or healthy CT training data New data
Figure 4: Flow chart showing process of fitting and use of the statistical shape model.
@@ 2.5 Testing and Comparison with Angle Based Classification
The testing and comparison were based on 92 data sets, which were not used for training of the SSMs and classifiers. These were from 40 control and 52 dysplasia subjects, for these subjects the current clinically accepted LCEA and AI angles were available. Angle-based classification was performed by combining the LCEA and AI angles, using LCEA (<20°)(<\$\$) and AI (>15°>\$\$) to indicate the presence of dysplasia. The comparison between the angle-based and two SSM-based classifers, namely LDDMM and GPDSSM, performance was made by plotting ROC curves and computing AUC scores.
2.6 Cross-validation
In order to formally test the hypothesis that our GPDSSM model outperformed the angle-based classification, we performed a leave-one-out cross-validation procedure (LOOCV) using a bootstrapping approach, resampled $100$ times over each fold to obtain a distribution of AUCs for both methods and estimate the test error. The LOOCV procedure also allowed the estimation of confidence intervals for accuracy, sensitivity and specificity.
We then performed a Wilcoxon signed rank test on the obtained distributions of AUCs. The null hypothesis posits there is no difference in the median AUC, and the alternate that the AUC distribution of the GPDSSM is stochastically greater than the angle AUC distribution. The test was performed on the difference distribution with $\alpha=0.05$.
@@ 3.2 Classification Performance
The GPDSSM outperforms the angle-based representation in terms of AUC (96.296.2% to 91.291.2%), suggesting that the overall classification based on our latent-space representation is more discriminative across all classification thresholds in separating dysplastics from controls. The GPDSSM representation also yields an improved sensitivity (performance on dysplastic) over the angle-based representations (94.294.2% to 84.684.6%), suggesting that this representation is particularly good at identifying dysplastic cases. In contrast, angle-based representations are more likely to incorrectly classify dysplastic as controls. This is an important area of improvement, as type II errors are far more costly in such a clinical setting, where patients may face significant deterioration of health if not treated due to false negatives.
In Table 1, the GPDSSM outperforms the standard LDDMM across all scores, which indicates the increased expressivity and powerful nonlinear dimensionality reduction afforded by our GPLVM-based diffeomorphic framework. We also remark that while the angle-based method has the best specificity (performance on controls) of 9090%, it underperforms in terms of sensitivity against both LDDMM (92.392.3% to 84.684.6%) and GPDSSM (94.294.2% to 84.684.6%). This suggests the use of latent-space representations based on the entire three-dimensional geometry has a measurable advantage in detecting dysplastic cases compared to the simpler angle-based representations.
@@ 4 Discussion
In this work, we have developed a semi-automated algorithm to classify the presence of dysplasia from volumetric CT scans, using a novel Gaussian process Statistical Shape Model. We have demonstrated the performance of our method on a dataset of real-world patient scans, our GPDSSM model can reliably detect the presence of dysplasia. Furthermore, our GPDSSM method (accuracy 90.2% and AUC 96.2%) demonstrably improves on predictions based on classical angle measures such as AI and LCEA (accuracy 87.0% and AUC 91.2%), as well as traditional LDDMM-based computational anatomy methods (accuracy 88.0% and AUC 92.7%). Our model utilises the full geometry of the acetabular surface to detect variations most correlated with dysplastic status, and is not dependent on a fixed 22D CT slice and angle of acquisition/measurement accuracy. The speed of inference and classification algorithm can save valuable time for clinicians, removing the need for manually marking and measuring angles, as well as improving detection accuracy. Given a new patient, clinicians may use our trained GPDSSM to infer dysplastic status as follows.
@@
The whole process, given surface data, takes a couple of minutes, though some manual monitoring is required to ensure the algorithm correctly picks out the acetabular correctly. If the patient is dysplastic, one may further use the type of deformation heat-map analysis presented in Section 2.7, for understanding the mode of disease variation in a patient scan, via the following steps.
@@ A.1 Model specification, fitting and inference
where we further condition on $p(U)$ a prior distribution over inducing variables $U$ [46] in time and latent space, where $U$ is equal to $\alpha$ evaluated at $I_{U}$, a set of $m<N$ free inducing points to be optimized in $0,1]\times Z$. One defines $p(\bm{\alpha}\mid U,Z)$ through standard Gaussian conditioning formulae.
@@ C.1 Gaussian Process Latent Variable Model
where $y$ is observed data, and $f:Z\longrightarrow Y$ is a randomized mapping from latent space Z{Z}, to Y{Y} the data space. The noise in the data is typically modelled by a Gaussian random variable. Importantly, $f$ is a Gaussian process [47], meaning the evaluations of $f$ at any finite set of latent points has a joint Gaussian distribution with mean and covariance determined by a mean function $m_{\theta}(z)$ and covariance kernel $k_{\theta}(z,z^{\prime})$ with learnable hyperparameters $\theta$. This is often denoted compactly as
@@ C.4 Classification of shape data
obtained and pre-processed with the methodology presented in Section 2, where $S_{i}$ denote the surface of the $i$’th subject anatomy and $y_{i}\in\{0,1\}$ the corresponding label; $1$ for dysplastics and 0 for controls. After fitting $\mathcal{M}$ to the shapes $S_{i}$, this yields a set